# On an approach to canonicalizing elliptic Feynman integrals

### v2

@@ 2.3 The transformation to canonical form
$\lim_{\lambda\to 0}{\bm{A}}_{\lambda}^{\rm can}=-\varepsilon\begin{pmatrix}\frac{\beta_{23}}{2\lambda}&0&\cdots&0&\frac{1}{\lambda}\\ 0&0&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&0&0\\ \frac{\beta_{23}^{2}}{4\lambda}&0&\cdots&0&\frac{\beta_{23}}{2\lambda}\end{pmatrix}\,,$ (40)
which is of UT and has only a simple pole. While the transformation above is only holomorphic around $\lambda=0$ but behaves logarithmically around $\lambda=1$, which is undesirable. This problem can be solved by simply changing $\varpi_{0}$ and $\vartheta_{0}$ into $\varpi_{1}$ and $\vartheta_{1}$ like Duhr:2024bzt thanks to the simple structure of the transformation ${\cal T}_{\rm can}$. Since the canonical basis here has a constant intersection matrix and thus is self-dual Pogel:2024sdi ; Duhr:2024xsy (up to a constant rotation) by construction, it seems close to the definition of the canonical basis advocated in Duhr:2024uid . Moreover, if one considers the restrictions in Dlapa:2022wdu , we should check the elements of connection matrices. Although we suppress explicit expressions for simplicity, we claim only $\varpi_{0/1}(\lambda)$ and no $\varpi^{\prime}_{0/1}(\lambda)$ appear (but we still have $\vartheta_{0/1}(e_{i},\lambda)$’s), with minimum degree of $\varpi_{0/1}(\lambda)$ equal to $-2$, satisfying the conditions in Dlapa:2022wdu .
It is worth discussing more about the $\varepsilon$-dependence in ${\cal T}_{\rm can}$. It is sometimes more favorable to work with a basis whose transformation to canonical basis (under the maximal cut) is independent of $\varepsilon$ and contains transcendental functions as less as possible, since it can help simplify the mixing with sub-sectors. Although ${\cal T}$ is independent of $\varepsilon$, it seems to contain more parts with transcendental functions than necessary, so we want to find a simpler basis. However, the $\varepsilon$-dependence cannot factorize out with elliptic integrals entirely, and they “entangle”. Fortunately, the entanglement only exists in the last row. More precisely, it only appears in the first and the last element in the last row of ${\cal T}_{\rm can}$. This implies that $\varepsilon$ only relates to $I_{n-1}$, which corresponds to the elliptic integrals of the second kind. Due to the simplicity, one could try disentangling the $\varepsilon$-dependent and the $\varepsilon$-independent parts in ${\cal T}_{\rm can}$. A possible decomposition is as follows:
${\cal T}_{\rm can}={\cal T}_{\rm can}^{(0)}\cdot{\cal T}_{\rm can}^{(\varepsilon)}\,,$ (41)
where
@@
${\cal T}_{\rm can}^{(\varepsilon)}=\begin{pmatrix}\frac{1}{\varpi_{0}(\lambda)}&0&\cdots&0&0\\ 0&1&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&1&0\\ -\frac{\lambda[\varpi_{0}(\lambda)-2(1-\lambda)\varpi_{0}^{\prime}(\lambda)]}{8\varepsilon}&0&\cdots&0&\frac{(1+2\beta_{1}\varepsilon)\varpi_{0}(\lambda)}{8\varepsilon}\end{pmatrix}\,,$ (42)
and then the form of ${\cal T}_{\rm can}^{(0)}$ is straightforward, which we suppress here. We see the elliptic integrals in ${\cal T}_{\rm can}^{(\varepsilon)}$ are elliptic integrals of the first and second kind, which are relatively simple to handle. Moreover, with ${\cal T}_{\rm can}^{(\varepsilon)}$, the $\varepsilon^{0}$-part of the connection matrices are strictly lower triangular. We also want to stress here though the transformation ${\cal T_{\rm can}}$ from pre-canonical basis $\vec{I}$ to canonical one depends on $\varepsilon$, the dependence is relatively simple and it is sometimes easier to simplify the mixing with sub-sectors in $\vec{I}$ because of the absence of elliptic integrals.
The above setup uses the Legendre form of the elliptic curve. We demonstrate how to transform the above pre-canonical basis back in terms of the general $u$ function in Eq. (3) in the following.
2.4 Transformation of the elliptic curve
For a general elliptic Feynman integral depending on multiple scales, its Baikov representation under the maximal cut is usually a one-fold integral, whose integrand is defined by the square root of a degree-$4$ polynomial $P_{4}(z)$, characterizing the elliptic curve. Here we suppress the twist part proportional to $\varepsilon$. We must stress that it is not always the case, there can be families defined with multivariate Baikov representations, even with the maximal cut, e.g. Delto:2023kqv . A possible solution is to construct the integrals in the form ${\rm d}\log\wedge{\rm d}\phi^{\rm E}$, where $\phi^{\rm E}$ denotes the elliptic part. However, it is beyond the scope of the paper and we will mainly focus on the univariate case.
It is well-known that Möbius transformations are automorphisms of the Riemann sphere, or equivalently, $\mathbb{CP}^{1}$. Two elliptic curves are isomorphic if they are related by a Möbius transformation. Leverage with Möbius transformation, one can always fix three out of the four branch points in $P_{4}(z)$ to $\{0,1,\infty\}$ and the other one to the modulus $\lambda$. In other words, one transforms the original form into the Legendre form of the elliptic curve. The modulus $\lambda$ is a function of kinematical variables, and thus it can be a function of several variables in general. Since a Möbius transformation is an automorphism of the Riemann sphere, the structure of the integral family remains and the complexity of kinematical dependence therein will not increase drastically. In the following, we demonstrate the general procedure for this transformation and then present the form of the pre-canonical basis written in terms of $P_{4}(z)$.
A Möbius transformation can be expressed as follows:
@@
The inverse transformation up to a rescaling is given by
@@
We now show how to transform an elliptic curve in a generic quartic form to the Legendre normal form. As mentioned above, Eq. (13), we require the transformation to map the branch points according to
@@
The form of transformation can then be easily deduced as
@@
With the relations above, we can now rewrite the pre-canonical basis in Eq. (18) in terms of the variable $z$. They are given by [^9] We leave out kinematic factors whose powers are proportional to $\varepsilon$, which would not break the $\varepsilon$-factorization.
$\phi_{1}$ $=\sqrt{c_{13}c_{24}}\frac{{\rm d}z}{\sqrt{P_{4}(z)}}\,,$ (52a)
$\phi_{i-3}$ $=\sqrt{P_{4}(c_{i})}\left(\frac{1}{z-c_{i}}-\frac{1}{c_{1i}}\right)\frac{{\rm d}z}{\sqrt{P_{4}(z)}}\,,\quad(i=5,\cdots,n)\,,$ (52b)
$\phi_{n-2}$ $=-(z-c_{1})\frac{{\rm d}z}{\sqrt{P_{4}(z)}}\,,$ (52c)
$\phi_{n-1}$ $=\frac{c_{41}}{c_{42}}\sqrt{c_{13}c_{24}}\frac{z-c_{2}}{z-c_{1}}\frac{{\rm d}z}{\sqrt{P_{4}(z)}}\,,$ (52d)
where we’ve assumed $c_{1}<c_{2}<c_{3}<c_{4}$.
It is interesting to study the asymptotic behaviors of the basis integrals in certain degenerate limits. Here we assume $c_{1}<c_{2}<z<c_{3}<c_{4}<c_{i}$ for concreteness where $i=5,\cdots,n$, and then $e_{i},e_{\infty}>1$. We consider the degenerate limit $c_{1}\to c_{2}$, where the genus of the curve drops. In this limit, the first two basis integrals asymptotically approach the same ${\rm d}\log$-form
$\lim_{c_{1}\to c_{2}}\phi_{1}=\lim_{c_{1}\to c_{2}}\phi_{n-1}=\frac{\sqrt{c_{13}c_{14}}}{(z-c_{1})\sqrt{(z-c_{3})(z-c_{4})}}{\rm d}z=-{\rm d}\log\left(\frac{1+\sqrt{\frac{c_{41}(z-c_{3})}{c_{31}(z-c_{4})}}}{1-\sqrt{\frac{c_{41}(z-c_{3})}{c_{31}(z-c_{4})}}}\right)\,.$ (53)
Similarly, $\phi_{i-3}$’s and $\phi_{n-1}$ approach two ${\rm d}\log$-forms as well:
$\lim_{c_{1}\to c_{2}}\phi_{i-3}$ $=\frac{\sqrt{c_{i3}c_{i4}}}{(z-c_{i})\sqrt{(z-c_{3})(z-c_{4})}}{\rm d}z={\rm d}\log\left(\frac{1+\sqrt{\frac{c_{i4}(z-c_{3})}{c_{i3}(z-c_{4})}}}{1-\sqrt{\frac{c_{i4}(z-c_{3})}{c_{i3}(z-c_{4})}}}\right)\,,$ (54)
$\lim_{c_{1}\to c_{2}}\phi_{n-2}$ $=-\frac{{\rm d}z}{\sqrt{(z-c_{3})(z-c_{4})}}={\rm d}\log\left(\frac{1+\sqrt{\frac{z-c_{3}}{z-c_{4}}}}{1-\sqrt{\frac{z-c_{3}}{z-c_{4}}}}\right)\,.$ (55)
Such properties are helpful if one wants to find “uniform-weight” master integrals in the elliptic sectors.
In order to obtain the canonical basis, one only needs to apply ${\cal T}_{\rm can}$ in Sec. 2.3, which is valid both for the Legendre family and the original family. The explicit expression for it can be obtained by plugging in
$\lambda=\frac{c_{14}c_{23}}{c_{13}c_{24}}\,,\quad e_{\infty}=\frac{c_{14}}{c_{24}}\,,\quad e_{i}=\frac{c_{14}c_{2i}}{c_{1i}c_{24}}\,,\;(i=5,\cdots,n)\,.$ (56)
It is straightforward to check the canonical basis integrals are also ${\rm d}\log$-forms asymptotically in limit $c_{1}\to c_{2}$, where the transformation only involves periods with subscripts “$1$” since $\lambda\to 1$ in this limit. We would like to mention that different choices of the transformation from the original family to the Legendre family as Eqs. (47) can result in a different pre-canonical basis with a different $\lambda$, and they are asymptotically ${\rm d}\log$-forms around different degenerate points, up to a constant transformation. This implies we can use different pre-canonical basis and their corresponding canonical basis in the neighborhood of different degenerate points to write the solutions of the integrals in the family, and finally cover the whole kinematic space, similar to Duhr:2024bzt . Although the pre-canonical basis in Eqs. (52) may not own the simplest asymptotic behaviors around the specific degenerate limit under consideration, for simplicity and concreteness, we will always use it for demonstration in the following.
@@ 3 Examples
In this section, we apply our results to examples to illustrate the above general approach. We consider the fully massive sunrise family and the non-planar double box family with an internal massive loop, and we include two different mass configurations for each family. For the non-planar double box family, the first configuration appears in di-jet and di-photon production Becchetti:2023wev , as well as Higgs boson pair production in small-mass limit Xu:2018eos ; Wang:2020nnr . It has also been discussed in earlier work Ahmed:2024tsg ; Becchetti:2025rrz . The other configuration we consider contributes to NNLO electroweak Møller scattering, and it is reported in another recent work Schwanemann:2024kbg .
3.1 Sunrise
3.1.1 Definition
Figure 1: The unequal-mass sunrise diagram is shown above, where different colors represent different masses for the propagators. The only elliptic sector in the family is the top sector. The sub-sectors are obtained by pinching the propagators, which are double tadpoles. We will first consider the two unequal masses case and then move to the case with three unequal masses. These cases are well-studied in the literature Bogner:2019lfa ; Gorges:2023zgv .
The unequal masses sunrise diagram is shown in Fig. 1. The external legs are massive, and $D_{1}$, $D_{2}$ and $D_{3}$ are three massive propagators with distinctive masses shown in Fig. 1, while $D_{4}$ and $D_{5}$ are two ISPs, i.e., $v_{4}\leq 0,v_{5}\leq 0$. They are defined as
$D_{1}$ $=k_{1}^{2}-m_{1}^{2}\,,\quad$ $D_{2}$ $=(k_{2}-p)^{2}-m_{2}^{2}\,,\quad$ $D_{3}$ $=(k_{1}-k_{2})^{2}-m_{3}^{2}\,,$ (57)
$D_{4}$ $=k_{2}^{2}\,,\quad$ $D_{5}$ $=(k_{1}-p)^{2}\,.$

### v3

@@ 2.3 The transformation to the canonical basis
$\lim_{\lambda\to 1}{\bm{A}}_{\lambda}^{\rm can}{\rm d}\lambda=-\varepsilon\begin{pmatrix}\frac{\beta_{34}}{2}&0&\cdots&0&1\\ 0&0&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&0&0\\ \frac{\beta_{34}^{2}}{4}&0&\cdots&0&\frac{\beta_{34}}{2}\end{pmatrix}{\rm d}\log(1-\lambda)\,$, (36)
and similar behaviors hold for ${\bm{A}}_{e_{i}}^{\rm can}{\rm d}e_{i}$ as well. These properties satisfy the conditions for a canonical basis as advocated in Dlapa:2022wdu . Moreover, our canonical basis has a constant intersection matrix and thus is self-dual Pogel:2024sdi ; Duhr:2024xsy (up to a constant rotation) by construction. Such properties are required in the definition of a canonical basis in Duhr:2024uid . The above discussions, together with those around Eq. (12), indicate that our canonical basis is indeed “canonical” in a broad sense.
At this point, we would like to discuss a bit about the sub-sector dependence of our canonical basis. Although our construction works under the maximal cut with no information about the sub-sectors, the simple form of our pre-canonical basis (21) allows us to naturally promote them to Feynman integrals $\vec{I}^{\text{pre}}$ without any cuts. It turns out that, in the differential equations of many examples, the sub-sector dependence of $\vec{I}^{\text{pre}}$ is rather simple, and can be easily converted to an $\varepsilon$-factorized form. However, this simplicity could be spoiled by the rotation matrix ${\cal T}_{\rm can}$, especially due to its non-trivial $\varepsilon$-dependence. Fortunately, the $\varepsilon$-dependence in ${\cal T}_{\rm can}$ only appears in the last row. It is possible to disentangle the $\varepsilon$-dependence by decomposing ${\cal T}_{\rm can}$ as
${\cal T}_{\rm can}={\cal T}_{\rm can}^{(0)}\,{\cal T}_{\rm can}^{(\varepsilon)}\,$, (37)
where ${\cal T}_{\rm can}^{(0)}$ is $\varepsilon$-independent and
@@
${\cal T}_{\rm can}^{(\varepsilon)}=\begin{pmatrix}\frac{1}{\varpi_{1}(\lambda)}&0&\cdots&0&0\\ 0&1&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&1&0\\ \frac{\lambda[\varpi_{1}(\lambda)-2(1-\lambda)\varpi_{1}^{\prime}(\lambda)]}{8\varepsilon}&0&\cdots&0&-\frac{(1+2\beta_{1}\varepsilon)\varpi_{1}(\lambda)}{8\varepsilon}\end{pmatrix}\,$. (38)
We can see that ${\cal T}_{\rm can}^{(\varepsilon)}$ only involves $\varpi_{1}$ and its derivative, which is relatively easy to handle. Moreover, the $\varepsilon^{0}$-part in the differential equations of ${\cal T}_{\rm can}^{(\varepsilon)}\vec{I}^{\text{pre}}$ is strictly lower triangular within the top-sector, similar to the case in Eq. (31). These properties are useful to simplify the sub-sector dependence, leading to a canonical basis for the full integral family. In fact, the simple $\varepsilon$-dependence of ${\cal T}_{\rm can}^{(\varepsilon)}$ often makes it easier to simplify part of sub-sector dependence directly from the pre-canonical basis. We will elaborate on this procedure in the next Section.
2.4 The Möbius transformation
As a final ingredient in our construction, we now briefly discuss the Möbius transformation (8), that takes any univariate elliptic integral defined by the degree-4 polynomial $P_{4}(z)$ in Eq. (3) into the Legendre normal form. This leads to our final result (10) for the canonical basis in a generic elliptic integral family.
It is well-known that Möbius transformations are automorphisms of the Riemann sphere, or equivalently, $\mathbb{CP}^{1}$. Two elliptic curves are isomorphic if they are related by a Möbius transformation. A Möbius transformation can be expressed as follows:
@@ 2.4 The Möbius transformation
The inverse transformation is given by
@@
For our purpose, we would like to map three out of the four branch points in $P_{4}(z)$ to $\{0,1,\infty\}$ and the other one to the modulus $\lambda$. For concreteness, we assume $c_{1}<c_{2}<c_{3}<c_{4}$, and require
@@
The above requirement uniquely fixes the form of the transformation as
@@
With the relations above, we can already rewrite the canonical basis $\mathcal{T}_{\text{can}}\phi^{\text{pre}}$ obtained in the previous subsection in terms of the variable $z$, leading to Eq. (10). On the other hand, it is interesting to investigate the behavior of the pre-canonical basis in Eq. (21) under this transformation. They are given by
$\phi_{1}^{\text{pre}}$ $=\sqrt{c_{13}c_{24}}\frac{{\rm d}z}{\sqrt{P_{4}(z)}}\,$, (48a)
$\phi_{i-3}^{\text{pre}}$ $=\sqrt{P_{4}(c_{i})}\left(\frac{1}{z-c_{i}}-\frac{1}{c_{1i}}\right)\frac{{\rm d}z}{\sqrt{P_{4}(z)}}\,,\quad(i=5,\cdots,n)\,$, (48b)
$\phi_{n-2}^{\text{pre}}$ $=-(z-c_{1})\frac{{\rm d}z}{\sqrt{P_{4}(z)}}\,$, (48c)
$\phi_{n-1}^{\text{pre}}$ $=\sqrt{c_{13}c_{24}}\frac{c_{41}(z-c_{2})}{c_{42}(z-c_{1})}\frac{{\rm d}z}{\sqrt{P_{4}(z)}}\,.$ (48d)
In Eq. (12), we have studied the asymptotic behavior of our canonical basis in the degenerate limit $c_{1}\to c_{2}$. Interestingly, our pre-canonical basis also exhibits the nice property of $\mathrm{d}\log$-form behavior in this limit. The first two basis integrals asymptotically approach the same ${\rm d}\log$-form
$\lim_{c_{1}\to c_{2}}\phi_{1}^{\text{pre}}=\lim_{c_{1}\to c_{2}}\phi_{n-1}^{\text{pre}}=\frac{\sqrt{c_{13}c_{14}}}{(z-c_{1})\sqrt{(z-c_{3})(z-c_{4})}}{\rm d}z=-{\rm d}\log\left(\frac{1+\sqrt{\frac{c_{41}(z-c_{3})}{c_{31}(z-c_{4})}}}{1-\sqrt{\frac{c_{41}(z-c_{3})}{c_{31}(z-c_{4})}}}\right)\,.$ (49)
Similarly, $\phi_{i-3}^{\text{pre}}$ and $\phi_{n-1}^{\text{pre}}$ approach two ${\rm d}\log$-forms as well:
$\lim_{c_{1}\to c_{2}}\phi_{i-3}^{\text{pre}}$ $=\frac{\sqrt{c_{i3}c_{i4}}}{(z-c_{i})\sqrt{(z-c_{3})(z-c_{4})}}{\rm d}z={\rm d}\log\left(\frac{1+\sqrt{\frac{c_{i4}(z-c_{3})}{c_{i3}(z-c_{4})}}}{1-\sqrt{\frac{c_{i4}(z-c_{3})}{c_{i3}(z-c_{4})}}}\right)\,$, (50)
$\lim_{c_{1}\to c_{2}}\phi_{n-2}^{\text{pre}}$ $=-\frac{{\rm d}z}{\sqrt{(z-c_{3})(z-c_{4})}}={\rm d}\log\left(\frac{1+\sqrt{\frac{z-c_{3}}{z-c_{4}}}}{1-\sqrt{\frac{z-c_{3}}{z-c_{4}}}}\right)\,.$ (51)
As a final remark, we emphasize that the construction in this Section has focused on the kinematic region around the degenerate point $c_{1}\to c_{2}$ and thus $\lambda\to 1$. If one is interested in the limit $c_{2}\to c_{3}$ and thus $\lambda\to 0$, a canonical basis can be easily constructed by replacing $\varpi_{1}$ by $\varpi_{0}$ in Eq. (10), and replacing $\vartheta_{1}$ by
which satisfies the differential equation
$\frac{\partial\vartheta_{0}(e_{i},\lambda)}{\partial e_{i}}=-\frac{(e_{i}-\lambda)\varpi_{0}(\lambda)+2\lambda(1-\lambda)\varpi_{0}^{\prime}(\lambda)}{2\sqrt{P_{L}(e_{i})}}\,.$ (53)
In the meantime, the Möbius transformation should be modified accordingly. The functions $\varpi_{0}$ and $\vartheta_{0}$ are holomorphic around $\lambda\to 0$. As a result, one can show that the canonical basis degenerates to $\mathrm{d}\log$-forms in that limit.
@@ 3 Examples
Family Graph Scales Sectors References
Unequal-mass sunrise (sr) [Uncaptioned image] 44 All ($7$) Bogner:2019lfa
An elliptic sub-sector of $2$-loop $t\bar{t}W$ production (ttW) [Uncaptioned image] 55 Top ($3$) New, see also Becchetti:2025qlu
${\cal T}_{3F}$ branch of $2$-loop $W$-pair production via light quark-antiquark annihilation (t3f) [Uncaptioned image] 44 All ($15$) New, see also He:2024iqg
Non-planar double box A (npdbA) [Uncaptioned image] 33 All ($36$) Becchetti:2025rrz
Non-planar double box B (npdbB) [Uncaptioned image] 33 Top ($4$) ++ Triangle ($3$) Schwanemann:2024kbg
Table 1: The examples we will consider are listed in the above table. The first column shows the integral family, with the abbreviated name used in this paper given in brackets. The second column contains the Feynman diagram for the family. Thick, colored lines represent propagators with non-zero invariant masses, and distinct colors represent different masses. The third column gives the number of kinematic scales; the number of independent dimensionless parameters is one less than this value. The fourth column indicates the sectors we consider, where the basis integrals are transformed into canonical form. The number in brackets is the number of master integrals. The number in the bracket is the number of the master integrals. The last column provides references for the canonical bases of the families. Two of these results are new to this work, and their relevant citations are also given.
In this section, we apply our results to various examples to demonstrate the generality of the canonical integrands. The sunrise families are prototypes of elliptic Feynman integrals, they are well-studied in Gorges:2023zgv ; Bogner:2019lfa . We will use the unequal-mass sunrise family as a pedagogical example. We then apply our methods to two phenomenologically interesting integral families, whose canonical bases were not reported in the literature. The first one is relevant to $t\bar{t}W$ production Becchetti:2025qlu , while the second one is relevant to $W$-pair production He:2024iqg . We present their canonical bases for the first time. In particular, for the $W$-pair production case, we derive the canonical basis for the full family without any cuts. This benefits from the simplicity of the sub-sector dependence of our pre-canonical basis.
Two further integral families are discussed as a demonstration of two-variable constructions under next-to-maximal cut. These are non-planar double box families with two different mass configurations. The first configuration appears in di-jet and di-photon production Becchetti:2023wev , as well as Higgs boson pair production in the small-mass limit Xu:2018eos ; Wang:2020nnr . It has also been discussed in earlier works Ahmed:2024tsg ; Becchetti:2025rrz . The second configuration is relevant to a subset of the NNLO electroweak corrections to Møller scattering ($e^{-}e^{-}\to e^{-}e^{-}$). This subset involves the exchange of a single type of massive gauge bosons ($W^{\pm}$ or $Z^{0}$), and was recently discussed in Schwanemann:2024kbg . For each family, we select a next-to-maximal cuts where two elliptic sectors survive. We show that the sub-sector dependence can be greatly simplified with such two-variable constructions.
For reference, the examples discussed in this section are summarized in Tab. 1.
3.1 Unequal-mass Sunrise
Figure 1: The unequal-mass sunrise diagram is shown above, where different colors represent different masses for the propagators. The only elliptic sector in the family is the top sector. The sub-sectors are obtained by pinching the propagators, which are double tadpoles. The case is well-studied in the literature Bogner:2019lfa .
The unequal-mass sunrise diagram is shown in Fig. 1. The external legs are massive, and $D_{1}$, $D_{2}$ and $D_{3}$ are three massive propagators with distinctive masses shown in Fig. 1, while $D_{4}$ and $D_{5}$ are two ISPs, i.e., $v_{4}\leq 0,v_{5}\leq 0$. They are defined as
$D_{1}$ $=k_{1}^{2}-m_{1}^{2}\,$$\quad$ $D_{2}$ $=(k_{2}-p)^{2}-m_{2}^{2}\,$$\quad$ $D_{3}$ $=(k_{1}-k_{2})^{2}-m_{3}^{2}\,$, (54)
$D_{4}$ $=k_{2}^{2}\,$$\quad$ $D_{5}$ $=(k_{1}-p)^{2}\,.$